# A Spectral Framework for Graph Neural Operators: Convergence Guarantees and Tradeoffs

### v2

@@ II-A Graphons
Graphons can be seen both as generative models for graphs and as limit objects of convergent graph sequences [12, 3]. Using a graphon as a generative model, we can construct an $n$-node graph $G_{n}$ from $\mathbf{W}$} by sampling points $u_{i}\in[0,1]$—associated with nodes $i\in[1,2,...,n]$—and assigning edge weight ${\mathbf{W}}(u_{i},u_{j})$ to edge $(i,j)$ to obtain a complete weighted graph; or, connecting edge $(i,j)$ with probability $p_{ij}={\mathbf{W}}(u_{i},u_{j})$ to obtain a stochastic unweighted graph. In either case, we say $G_{n}$ is sampled from $\mathbf{W}$}.
@@ II-B Cut norm and cut distance
It can be shown that convergent graph sequences in the homomorphism density sense also converge in the cut distance. This statement is made formal by defining the graphons ${\mathbf{W}}_{G_{n}}$ induced by the graphs $G_{n}$, which allows comparison of the graphs $G_{n}$ and the graphon $\mathbf{W}$} through the above-defined distance. For an $n$-node graph $G$, such induced graphon is created by constructing a regular partition $I_{1}\cup...\cup I_{n}$ of $[0,1]$, where $I_{j}=[(j-1)/n,j/n],1\leq j\leq n-1$ and $I_{n}=[(n-1)/n,1]$, and defining
${\mathbf{W}}_{G}(u,v)=\sum_{j=1}^{n}\sum_{k=1}^{n}[A]_{jk}\mathbb{I}(u\in I_{j})\mathbb{I}(v\in I_{k}).$
@@ III-B Lipschitz case
Weyl’s inequality gives $|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|\leq\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}$. Since $\|T_{\mathbf{K}}\|_{2}\leq\|{\mathbf{K}}\|_{2}$, Lemma 4 directly implies $\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}\leq\|{\mathbf{W}}_{G_{n}}-{\mathbf{W}}\|_{2}$, which yields the stated rate $\mathcal{O}(\sqrt{\log(n)/n})$. ∎
@@ III-D Discussion
GNN transferability bounds typically decompose into two factors: a constant term that depends on structural and architectural assumptions (such as GNN depth), and a rate term that decays asymptotically with $n$, determined by spectral convergence. The focus of this note is on the operator-level rates, which can be directly substituted into any existing transferability bound, e.g., [13], to yield guarantees under different structural assumptions on the underlying graphon.
@@ IV Numerical examples
The synthetic graphon is generated from a smooth function over $0,1]^{2}$ to define connection probabilities. For Cora, we use the full adjacency matrix as the graphon, while for PubMed we use the adjacency matrix of a randomly sampled subset. We can normalize the graphon adjacency matrices, via the $L_{1}$ norm, and reorder by node degree. Leveraging polynomial interpolation we determine an approximate Lipschitz constant for the graphons and particular graphon “pieces”. Sorting and interpolating is consistent with the sort-and-smooth approach used in graphon estimation literature [4].

### v3

@@ II-A Graphons
Graphons, symmetric measurable functions ${\mathbf{W}}:[0,1]^{2}\rightarrow[0,1]$, can be seen both as generative models for graphs and as limit objects of convergent graph sequences [12, 3]. Using a graphon as a generative model, we can construct an $n$-node graph $G_{n}$ from $\mathbf{W}$} by sampling points $u_{i}\in[0,1]$—associated with nodes $i\in[1,2,...,n]$—and assigning edge weight ${\mathbf{W}}(u_{i},u_{j})$ to edge $(i,j)$ to obtain a complete weighted graph; or, connecting edge $(i,j)$ with probability $p_{ij}={\mathbf{W}}(u_{i},u_{j})$ to obtain a stochastic unweighted graph. In either case, we say $G_{n}$ is sampled from $\mathbf{W}$}.
@@ II-B Cut norm and cut distance
Convergent graph sequences in the homomorphism density sense also converge in the cut distance. This is formalized by defining graphons ${\mathbf{W}}_{G_{n}}$ induced by the graphs $G_{n}$, which allows comparison of graphs to graphons through the above-defined distance. For an $n$-node graph $G_{n}$, an induced graphon is created by constructing a regular partition $I_{1}\cup...\cup I_{n}$ of $[0,1]$, where $I_{j}=[(j-1)/n,j/n),1\leq j\leq n-1$ and $I_{n}=[(n-1)/n,1]$, and with $\mathbb{I}$ as the indicator function defining
${\mathbf{W}}_{G_{n}}(u,v)=\sum_{j=1}^{n}\sum_{k=1}^{n}[A]_{jk}\mathbb{I}(u\in I_{j})\mathbb{I}(v\in I_{k}).$
@@ III-B Lipschitz case
Weyl’s inequality gives $|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|\leq\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}$. Since $\|T_{\mathbf{K}}\|_{2}\leq\|{\mathbf{K}}\|_{2}$, $\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}\leq\|{\mathbf{W}}_{G_{n}}-{\mathbf{W}}\|_{2}$, and Lemma 4 directly implies the stated rate $\mathcal{O}(\sqrt{\log(n)/n})$. ∎
@@ III-D Discussion
GNN transferability bounds typically decompose into two factors: a constant term that depends on structural and architectural assumptions (such as GNN depth), and a rate term that decays asymptotically with $n$, determined by spectral convergence. The focus of this work is on the operator-level rates, which can be directly substituted into any existing transferability bound, e.g., [13], to yield guarantees under different structural assumptions on the underlying graphon.
@@ IV Numerical examples
The synthetic graphon is generated from a smooth function over $0,1]^{2}$ to define connection probabilities. For Cora, we use the full adjacency matrix as the graphon, while for PubMed we use the adjacency matrix of a randomly sampled subset. We normalize the graphon adjacency matrices, via the $L_{1}$ norm, and reorder by node degree. Leveraging polynomial interpolation we determine an approximate Lipschitz constant for the graphons and particular graphon “pieces”. Sorting and interpolating is consistent with the sort-and-smooth approach used in graphon estimation literature [4].